Question 2 (of 3) e: 0 points 5 out of 5 attem What is the smallest energy photon that can be absorbed by a hydrogen atom in its ninth excited state? eV References eBook & Resources Multipart Answer
Answer #1

For smallest energy absorption the photon must be able to excite electron to the next shell n1 = 10 n2 = 11

E = 13.6*(1/n1^2 - 1/n2^2) =  13.6*(1/10^2 - 1/11^2) = 0.0236 eV